OpenStudy (anonymous):

What is the slope of the line defined by the equation 8x+2y=5?

OpenStudy (icalibear):

8

OpenStudy (abb0t):

rearrange using algebra your slope should be 4

OpenStudy (icalibear):

once you divide by 2 its 4

OpenStudy (abb0t):

re-write it in y=mx+b

OpenStudy (anonymous):

You guys confused me... @icalibear @abb0t

OpenStudy (icalibear):

sorry the slope should be 4 since y = mx + b m = 8/2 = 4

OpenStudy (anonymous):

^ try again

OpenStudy (abb0t):

\[8x + 2y=5\] \[2y=5-8x\] \[y=\frac{ 5 }{ 2 }-\frac{ 8 }{ 2 }x\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

slope = gradient = (-4)